OpenStudy (anonymous):

Select all of the following that are quadratic equations. 3x^2 + 5x - 7 = 0 x^3 - 2x^2 + 1 = 0 2x - 1 = 0 5x^2+ 15x = 0 6x - 1 = 4x + 7 x^2 - 4x = 4x + 7

OpenStudy (anonymous):

3x^2 + 5x - 7 = 0 5x^2+ 15x = 0 x^2 - 4x = 4x + 7